# Matryoshka Quantization

### v1

@@ Appendix F Detailed Evaluation for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
int8 $[8,4,8\rightarrow 2]$ $57.59$ $77.27$ $81.83$ $75.48$ $81.01$ $67.25$ $73.4$ $2.467$
int4 $[8,4,8\rightarrow 2]$ $57.68$ $78.45$ $82.97$ $75.5$ $80.85$ $67.56$ $73.84$ $2.488$
int2 $[8,4,8\rightarrow 2]$ $48.81$ $74.03$ $81.65$ $68.1$ $77.48$ $65.11$ $69.2$ $2.796$
int6 $[8,4,8\rightarrow 2]$ $57.42$ $77.19$ $81.87$ $75.42$ $81.01$ $67.8$ $73.45$ $2.468$
int3 $[8,4,8\rightarrow 2]$ $55.63$ $75.88$ $80.12$ $74.01$ $80.36$ $67.96$ $72.33$ $2.549$
$[8,4,2,8\rightarrow 2]$ $54.35$ $76.85$ $79.33$ $74.6$ $80.47$ $67.4$ $72.17$ $2.543$
@@
int8 $[8,4,8\rightarrow 2]$ $58.11$ $76.43$ $81.25$ $79.12$ $82.05$ $71.35$ $74.72$ $2.298$
$[8,4,2,8\rightarrow 2]$ $57.51$ $76.43$ $81.53$ $78.95$ $82.1$ $71.19$ $74.62$ $2.299$
$[8,4,2,8\rightarrow 4;2]$ $58.11$ $76.14$ $81.68$ $79.12$ $82.26$ $71.51$ $74.8$ $2.302$
int4 $[8,4,8\rightarrow 2]$ $57.42$ $76.35$ $77.55$ $78.06$ $81.61$ $71.59$ $73.76$ $2.328$
$[8,4,2,8\rightarrow 2]$ $56.91$ $75.8$ $78.44$ $77.76$ $81.39$ $72.38$ $73.78$ $2.329$
$[8,4,2,8\rightarrow 4;2]$ $57.51$ $75.76$ $75.96$ $77.96$ $81.72$ $71.98$ $73.48$ $2.33$
int2 $[8,4,8\rightarrow 2]$ $39.51$ $65.03$ $66.88$ $63.37$ $75.08$ $61.01$ $61.81$ $2.74$
$[8,4,2,8\rightarrow 2]$ $40.78$ $66.5$ $67.55$ $63.67$ $75.95$ $60.62$ $62.51$ $2.746$
$[8,4,2,8\rightarrow 4;2]$ $40.19$ $65.7$ $65.57$ $63.83$ $75.3$ $62.12$ $62.12$ $2.746$
int6 $[8,4,8\rightarrow 2]$ $57.85$ $76.09$ $81.47$ $78.98$ $81.88$ $71.27$ $74.59$ $2.301$
$[8,4,2,8\rightarrow 2]$ $57.17$ $75.97$ $82.2$ $79$ $81.83$ $71.9$ $74.68$ $2.302$
$[8,4,2,8\rightarrow 4;2]$ $57.42$ $76.09$ $82.29$ $78.95$ $82.10$ $71.27$ $74.69$ $2.305$
int3 $[8,4,8\rightarrow 2]$ $51.96$ $71.55$ $78.07$ $73.17$ $79.43$ $66.93$ $70.18$ $2.485$
$[8,4,2,8\rightarrow 2]$ $50.94$ $71.76$ $78.78$ $73.09$ $79.05$ $66.77$ $70.06$ $2.486$
$[8,4,2,8\rightarrow 4;2]$ $51.45$ $72.39$ $78.84$ $73.46$ $79.6$ $67.96$ $70.62$ $2.731$
@@
S.P. ${\rm MatQuant}$ $24.91$ $41.62$ $62.26$ $40.87$ $63.38$ $53.67$ $47.78$
${\rm MatQuant}$ $28.24$ $39.23$ $62.17$ $39.13$ $63.49$ $50.75$ $47.17$
int6 Sliced int8 $58.53$ $77.15$ $82.48$ $79.04$ $81.5$ $68.67$ $74.56$
@@
S.P. ${\rm MatQuant}$ $23.21$ $23.82$ $37.83$ $24.67$ $49.02$ $49.57$ $34.69$
${\rm MatQuant}$ $22.27$ $32.49$ $62.02$ $32.43$ $59.3$ $51.46$ $43.33$
int6 Sliced int8 $49.32$ $73.53$ $82.66$ $80.16$ $81.12$ $75.45$ $73.71$
@@
${\rm MatQuant}$ $28.24$ $51.73$ $64.19$ $46.76$ $68.66$ $55.01$ $52.43$ $3.153$
Table 21: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$ and the Baseline for int2, int4, int8 quatization of Gemma-2 9B with OmniQuant. Note that the model was trained with Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ for int2, the int4 and int8 model were sliced post training.
@@
int8 S.P. ${\rm MatQuant}$ $56.48$ $76.85$ $73.36$ $74.87$ $80.74$ $66.77$ $71.51$ $2.525$
OmniQuant $59.47$ $77.31$ $83.94$ $77.35$ $81.39$ $68.11$ $74.59$ $2.418$
MatQuant $58.11$ $78.03$ $83.27$ $76.17$ $81.18$ $67.09$ $73.97$ $2.451$
int4 S.P. ${\rm MatQuant}$ $57.17$ $77.02$ $74.28$ $74.41$ $80.69$ $67.56$ $71.85$ $2.543$
OmniQuant $58.79$ $78.37$ $83.55$ $76.71$ $81.45$ $67.09$ $74.33$ $2.451$
MatQuant $57.25$ $77.36$ $84.86$ $75.52$ $81.5$ $66.77$ $73.88$ $2.481$
int2 S.P. ${\rm MatQuant}$ $49.74$ $74.66$ $80.92$ $66.57$ $76.06$ $63.54$ $68.58$ $2.857$
OmniQuant $39.16$ $63.43$ $72.11$ $52.24$ $72.63$ $61.88$ $60.24$ $3.292$
MatQuant $48.72$ $72.18$ $79.2$ $68.11$ $76.17$ $66.77$ $68.52$ $2.809$
Table 22: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$ and the Baseline for int2, int4, int8 quatization of Gemma-2 9B with QAT. Note that the model was trained with Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ for int2, the int4 and int8 model were sliced post training.
@@
int8 S.P. ${\rm MatQuant}$ $55.97$ $76.18$ $80.09$ $75.43$ $80.69$ $68.9$ $72.88$ $2.429$
QAT $47.78$ $70.66$ $75.08$ $69.92$ $78.35$ $65.11$ $67.82$ $2.29$
MatQuant $46.25$ $71.21$ $75.6$ $69.97$ $78.4$ $64.64$ $67.68$ $2.301$
int4 S.P. ${\rm MatQuant}$ $55.2$ $76.01$ $74.74$ $74.19$ $80.41$ $68.9$ $71.57$ $2.429$
QAT $46.16$ $71.59$ $73.73$ $68.72$ $78.62$ $63.38$ $67.03$ $2.324$
MatQuant $44.37$ $70.45$ $75.81$ $68.43$ $78.35$ $64.88$ $67.05$ $2.332$
int2 S.P. ${\rm MatQuant}$ $41.21$ $66.2$ $65.02$ $64.31$ $76.06$ $62.35$ $62.53$ $2.706$
QAT $33.45$ $55.43$ $62.26$ $54.8$ $70.51$ $59.67$ $56.02$ $2.923$
MatQuant $39.85$ $65.66$ $65.93$ $64.08$ $75.68$ $62.75$ $62.32$ $2.756$
Table 23: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$ and the Baseline for int2 quatization of Mistral 7B with OmniQuant and QAT.
@@
${\rm MatQuant}$ $34.3$ $55.09$ $71.83$ $65.89$ $75.52$ $65.11$ $61.29$ $2.474$

### v2

@@ Appendix H Detailed Evaluation for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$
int8 $[8,4,8\rightarrow 2]$ $57.51$ $76.26$ $83.30$ $73.35$ $80.74$ $65.43$ $72.76$ $2.473$
$[8,4,2,8\rightarrow 2]$ $58.19$ $76.89$ $83.73$ $76.75$ $81.39$ $67.01$ $73.99$ $2.435$
$[8,4,2,8\rightarrow 4;2]$ $57.68$ $77.06$ $83.00$ $76.76$ $81.45$ $67.17$ $73.85$ $2.437$
int4 $[8,4,8\rightarrow 2]$ $56.23$ $76.47$ $82.63$ $73.03$ $80.69$ $66.85$ $72.65$ $2.519$
$[8,4,2,8\rightarrow 2]$ $57.51$ $76.73$ $83.36$ $76.23$ $80.85$ $67.09$ $73.63$ $2.486$
$[8,4,2,8\rightarrow 4;2]$ $57.51$ $76.68$ $83.27$ $75.85$ $81.61$ $66.38$ $73.55$ $2.478$
int2 $[8,4,8\rightarrow 2]$ $38.14$ $66.50$ $76.73$ $59.70$ $71.11$ $63.69$ $62.64$ $3.289$
$[8,4,2,8\rightarrow 2]$ $40.61$ $67.55$ $71.07$ $60.80$ $72.96$ $64.48$ $62.91$ $3.138$
$[8,4,2,8\rightarrow 4;2]$ $42.75$ $69.65$ $74.40$ $60.53$ $72.42$ $66.14$ $64.32$ $3.227$
int6 $[8,4,8\rightarrow 2]$ $57.59$ $76.30$ $83.55$ $73.41$ $80.85$ $65.51$ $72.87$ $2.469$
$[8,4,2,8\rightarrow 2]$ $58.28$ $76.85$ $83.43$ $76.91$ $81.18$ $67.01$ $73.94$ $2.438$
$[8,4,2,8\rightarrow 4;2]$ $58.11$ $76.98$ $83.33$ $76.70$ $81.45$ $67.48$ $74.01$ $2.439$
int3 $[8,4,8\rightarrow 2]$ $52.30$ $75.25$ $78.26$ $71.08$ $79.49$ $65.35$ $70.29$ $2.651$
$[8,4,2,8\rightarrow 2]$ $54.44$ $75.97$ $82.20$ $73.84$ $80.20$ $66.46$ $72.19$ $2.603$
$[8,4,2,8\rightarrow 4;2]$ $54.44$ $76.26$ $81.90$ $73.89$ $79.92$ $65.75$ $72.03$ $2.604$
@@
int8 $[8,4,8\rightarrow 2]$ $57.68$ $76.09$ $82.60$ $78.75$ $82.48$ $70.88$ $74.75$ $2.242$
$[8,4,2,8\rightarrow 2]$ $57.76$ $76.35$ $81.50$ $79.13$ $82.43$ $72.06$ $74.87$ $2.240$
$[8,4,2,8\rightarrow 4;2]$ $58.19$ $76.05$ $81.62$ $78.92$ $82.21$ $71.90$ $74.81$ $2.240$
int4 $[8,4,8\rightarrow 2]$ $57.85$ $76.81$ $78.47$ $77.62$ $80.96$ $70.88$ $73.76$ $2.279$
$[8,4,2,8\rightarrow 2]$ $57.08$ $75.88$ $78.47$ $77.65$ $81.34$ $72.22$ $73.77$ $2.276$
$[8,4,2,8\rightarrow 4;2]$ $57.34$ $75.80$ $78.99$ $77.67$ $81.50$ $72.30$ $73.93$ $2.277$
int2 $[8,4,8\rightarrow 2]$ $40.61$ $67.17$ $67.37$ $63.10$ $75.24$ $60.38$ $62.31$ $2.670$
$[8,4,2,8\rightarrow 2]$ $40.53$ $66.71$ $67.89$ $63.29$ $75.46$ $62.35$ $62.70$ $2.673$
$[8,4,2,8\rightarrow 4;2]$ $40.10$ $66.37$ $67.86$ $63.14$ $75.08$ $63.06$ $62.60$ $2.670$
int6 $[8,4,8\rightarrow 2]$ $57.85$ $76.05$ $82.23$ $78.70$ $82.10$ $71.43$ $74.73$ $2.245$
$[8,4,2,8\rightarrow 2]$ $58.11$ $75.93$ $82.14$ $79.10$ $82.26$ $71.19$ $74.79$ $2.243$
$[8,4,2,8\rightarrow 4;2]$ $58.19$ $75.67$ $81.31$ $78.80$ $82.15$ $71.27$ $74.56$ $2.243$
int3 $[8,4,8\rightarrow 2]$ $51.19$ $71.00$ $76.67$ $73.07$ $79.54$ $68.03$ $69.92$ $2.441$
$[8,4,2,8\rightarrow 2]$ $51.71$ $71.46$ $76.85$ $73.00$ $79.00$ $67.88$ $69.98$ $2.437$
$[8,4,2,8\rightarrow 4;2]$ $51.28$ $71.34$ $76.12$ $72.96$ $79.33$ $68.98$ $70.00$ $2.435$
@@
S.P. ${\rm MatQuant}$ $25.26$ $38.47$ $62.14$ $35.09$ $61.70$ $51.46$ $45.69$
${\rm MatQuant}$ $23.72$ $36.62$ $62.17$ $33.72$ $59.36$ $49.57$ $44.19$
int6 Sliced int8 $58.53$ $77.10$ $83.00$ $78.81$ $81.07$ $68.43$ $74.49$
@@
S.P. ${\rm MatQuant}$ $22.53$ $25.51$ $38.90$ $24.13$ $50.92$ $50.12$ $35.35$
${\rm MatQuant}$ $21.33$ $25.59$ $57.37$ $24.85$ $50.92$ $50.12$ $38.36$
int6 Sliced int8 $49.32$ $73.53$ $82.60$ $80.28$ $80.96$ $74.98$ $73.61$
@@
${\rm MatQuant}$ $28.33$ $51.85$ $63.64$ $46.94$ $68.28$ $54.14$ $52.20$ $3.055$
Table 23: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$ and the Baseline for int2, int4, int8 quatization of Gemma-2 9B with Baseline. Note that the model was trained with Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ for int2; the int4 and int8 model were sliced post training.
@@
int8 S.P. ${\rm MatQuant}$ $57.94$ $76.64$ $82.66$ $76.98$ $81.01$ $67.56$ $73.80$ $2.372$
Baseline $59.47$ $77.31$ $83.94$ $77.35$ $81.39$ $68.11$ $74.59$ $2.418$
${\rm MatQuant}$ $57.59$ $77.02$ $84.01$ $76.61$ $81.18$ $67.88$ $74.05$ $2.438$
int4 S.P. ${\rm MatQuant}$ $57.17$ $76.39$ $81.47$ $75.81$ $80.85$ $66.38$ $73.01$ $2.420$
Baseline $58.79$ $78.37$ $83.55$ $76.71$ $81.45$ $67.09$ $74.33$ $2.451$
${\rm MatQuant}$ $58.02$ $78.11$ $83.24$ $76.08$ $80.96$ $66.54$ $73.83$ $2.491$
int2 S.P. ${\rm MatQuant}$ $40.44$ $66.75$ $77.92$ $60.42$ $72.52$ $66.06$ $64.02$ $3.171$
Baseline $39.16$ $63.43$ $72.11$ $52.24$ $72.63$ $61.88$ $60.24$ $3.292$
${\rm MatQuant}$ $40.78$ $67.85$ $73.64$ $60.56$ $72.09$ $65.19$ $63.35$ $3.187$
Table 24: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$ and the Baseline for int2, int4, int8 quatization of Gemma-2 9B with Baseline. Note that the model was trained with Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ for int2; the int4 and int8 model were sliced post training.
@@
int8 S.P. ${\rm MatQuant}$ $55.89$ $75.84$ $79.57$ $75.47$ $81.07$ $68.43$ $72.71$ $2.363$
Baseline $58.11$ $75.38$ $80.12$ $78.7$ $81.5$ $71.19$ $74.17$ $2.29$
MatQuant $57.68$ $76.09$ $82.23$ $78.41$ $82.26$ $70.48$ $74.52$ $2.262$
int4 S.P. ${\rm MatQuant}$ $54.95$ $75.59$ $75.05$ $74.60$ $80.79$ $69.06$ $71.67$ $2.394$
Baseline $56.91$ $75.42$ $75.38$ $78.06$ $81.39$ $72.38$ $73.26$ $2.324$
MatQuant $56.66$ $75.72$ $77.55$ $77.30$ $81.23$ $70.96$ $73.24$ $2.295$
int2 S.P. ${\rm MatQuant}$ $40.53$ $67.38$ $66.91$ $63.62$ $75.63$ $61.88$ $62.66$ $2.656$
Baseline $33.45$ $55.43$ $62.26$ $54.8$ $70.51$ $59.67$ $56.02$ $2.923$
MatQuant $41.21$ $66.84$ $65.41$ $63.61$ $75.41$ $61.25$ $62.29$ $2.660$
Table 25: Table presents downstream evaluation and perplexity results for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, comparing it with ${\rm MatQuant}$, and the Baseline for int2 quatization of Mistral 7B. Results are presented for both, OmniQuant and QAT as the base algorithms.
@@
${\rm MatQuant}$ $35.58$ $56.36$ $72.66$ $66.68$ $74.32$ $66.22$ $61.97$ $2.524$
Table 26: Table presents the downstream evaluation results for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ when applied to OmniQuant on Gemma-2 2B.
Avg. Bits Method Gemma-2 2B
OmniQuant ARC-c ARC-e BoolQ HellaSwag PIQA Winogrande Average
bfloat16 $50.09$ $71.59$ $76.45$ $69.69$ $78.29$ $63.14$ $68.21$
8 ${\rm MatQuant}$ $49.66$ $71.00$ $76.73$ $68.85$ $78.56$ $63.30$ $68.02$
8 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $48.04$ $71.8$ $75.78$ $67.64$ $78.07$ $63.22$ $67.42$
4 ${\rm MatQuant}$ $47.27$ $70.79$ $73.76$ $66.85$ $78.07$ $62.75$ $66.58$
4.023 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $45.65$ $70.29$ $74.8$ $66.07$ $77.58$ $62.27$ $66.11$
2 ${\rm MatQuant}$ $29.95$ $54.21$ $64.40$ $44.37$ $66.81$ $54.46$ $52.37$
2.052 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $34.39$ $59.64$ $62.69$ $52.11$ $69.86$ $55.56$ $55.71$
6 ${\rm MatQuant}$ $48.89$ $70.50$ $75.69$ $68.89$ $78.40$ $62.75$ $67.52$
6.018 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $47.1$ $71.46$ $76.02$ $67.47$ $77.91$ $63.61$ $67.26$
3 ${\rm MatQuant}$ $44.03$ $67.09$ $74.25$ $62.78$ $77.26$ $61.40$ $64.47$
3.031 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $44.45$ $68.56$ $69.11$ $62.28$ $75.95$ $62.59$ $63.82$
Table 27: Table presents the downstream evaluation results for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ when applied to OmniQuant on Gemma-2 9B.
Avg. Bits Method Gemma-2 9B
OmniQuant ARC-c ARC-e BoolQ HellaSwag PIQA Winogrande Average
bfloat16 $58.96$ $77.57$ $83.33$ $77.31$ $81.12$ $67.96$ $74.38$
8 ${\rm MatQuant}$ $57.59$ $77.02$ $84.01$ $76.61$ $81.18$ $67.88$ $74.05$
8 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $58.11$ $78.03$ $83.27$ $76.17$ $81.18$ $67.09$ $73.97$
4 ${\rm MatQuant}$ $58.02$ $78.11$ $83.24$ $76.08$ $80.96$ $66.54$ $73.83$
4.022 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $57.25$ $77.36$ $84.86$ $75.52$ $81.5$ $66.77$ $73.88$
2 ${\rm MatQuant}$ $40.78$ $67.85$ $73.64$ $60.56$ $72.09$ $65.19$ $63.35$
2.050 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $48.72$ $72.18$ $79.2$ $68.11$ $76.17$ $66.77$ $68.52$
6 ${\rm MatQuant}$ $57.25$ $76.94$ $84.04$ $76.63$ $81.34$ $67.32$ $73.92$
6.018 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $58.87$ $78.03$ $83.61$ $76.18$ $81.45$ $67.09$ $74.21$
3 ${\rm MatQuant}$ $55.80$ $76.89$ $81.99$ $74.27$ $80.14$ $68.11$ $72.87$
3.029 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $55.46$ $76.14$ $84.04$ $74.49$ $80.14$ $67.32$ $72.93$
Table 28: Table presents the downstream evaluation results for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ when applied to OmniQuant on Mistral 7B.
Data type Method Mistral 7B
OmniQuant ARC-c ARC-e BoolQ HellaSwag PIQA Winogrande Average
bfloat16 $49.57$ $73.74$ $84.4$ $80.61$ $81.18$ $74.43$ $73.99$
8 ${\rm MatQuant}$ $49.06$ $72.52$ $84.74$ $79.21$ $81.45$ $74.90$ $73.65$
8 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $48.04$ $73.44$ $84.13$ $79.37$ $81.12$ $74.66$ $73.46$
4 ${\rm MatQuant}$ $47.87$ $71.55$ $83.88$ $78.85$ $81.34$ $74.90$ $73.06$
4.022 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $48.21$ $72.69$ $83.49$ $78.82$ $81.12$ $74.43$ $73.13$
2 ${\rm MatQuant}$ $37.88$ $62.58$ $73.15$ $65.89$ $73.88$ $63.14$ $62.75$
2.051 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $41.38$ $67.42$ $71.62$ $71.98$ $77.86$ $65.67$ $65.99$
6 ${\rm MatQuant}$ $49.40$ $72.47$ $84.68$ $79.52$ $81.34$ $74.35$ $73.63$
6.018 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $48.46$ $72.98$ $84.07$ $79.64$ $81.18$ $75.22$ $73.59$
3 ${\rm MatQuant}$ $47.35$ $71.00$ $80.00$ $76.96$ $80.30$ $71.35$ $71.16$
3.030 Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $45.65$ $71.21$ $80.43$ $78.31$ $81.07$ $72.61$ $71.55$
Table 29: Table presents the downstream evaluation and perplexity results for our Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ co-distillation experiments on Gemma-2 9B with OmniQuant.
OmniQuant Gemma-2 9B
Avg. Bits Config. ARC-c ARC-e BoolQ HellaSwag PIQA Winogrande Average log pplx.
$8$ $[8,4,8\rightarrow 2]$ $57.59$ $77.27$ $81.83$ $75.48$ $81.01$ $67.25$ $73.4$ $2.467$
$4.022$ $[8,4,8\rightarrow 2]$ $57.68$ $78.45$ $82.97$ $75.5$ $80.85$ $67.56$ $73.84$ $2.488$
$2.050$ $[8,4,8\rightarrow 2]$ $48.81$ $74.03$ $81.65$ $68.1$ $77.48$ $65.11$ $69.2$ $2.796$
6.018 $[8,4,8\rightarrow 2]$ $57.42$ $77.19$ $81.87$ $75.42$ $81.01$ $67.8$ $73.45$ $2.468$
$3.029$ $[8,4,8\rightarrow 2]$ $55.63$ $75.88$ $80.12$ $74.01$ $80.36$ $67.96$ $72.33$ $2.549$
$[8,4,2,8\rightarrow 2]$ $54.35$ $76.85$ $79.33$ $74.6$ $80.47$ $67.4$ $72.17$ $2.543$